An ordinary (fair) die is a cube with the numbers through on the sides (represented by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This sum is recorded as the outcome of a single trial of a random experiment.
Compute the probability of each of the following events:
Event : The sum is greater than 7 .
Event : The sum is not divisible by 6.
Write your answers as exact fractions.


Answer by ewatrrr(24785)


Hi,
note: the '36' possible outcomes of sums rolling 2 die
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12
P(The sum > 7 ) = 15/36 or 5/12
P(The sum is not divisible by 6) = 1 - 6/36 -1/36 = 29/36